OpenStudy (anonymous):

The sum of the squares of two consecutive positive even integers is 340. Find the integers. @noemy_bieber

OpenStudy (anonymous):

let x be the first number then x + 2 is the next consecutive even integer then x^2 + (x + 2)^2 = 340 x^2 + x^2 + 4x + 4 = 340 2x^2 + 4x - 336 = 0 x^2 +2x - 168 = 0 (x + 14)(x - 12) = 0 x = -14 or 12 so the two numbers are 12 and 14 since they need to be positive

OpenStudy (anonymous):

x=12, x+2=14